OpenStudy (anonymous):

Verify the identity. (1-sin x /cos x )= (cos x/ 1+ sin x)

OpenStudy (anonymous):

its just Pythagorean identity \[\sin ^{2}\left( x \right)+\cos ^{2}\left( x \right)=1\]

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

\[\cos ^{2}\left( x \right)=1-\sin ^{2}\left( x \right)\]\[\cos\left( x \right)\cos \left( x \right)=\left( 1-\sin \left( x \right) \right)\left( 1+\sin \left( x \right) \right)\]\[\frac{ \cos \left( x \right) }{ 1-\sin \left( x \right) }=\frac{ 1+\sin \left( x \right) }{ \cos \left( x \right) }\]

OpenStudy (anonymous):

is that an answer?

OpenStudy (anonymous):

yes